a.      When the price of mayonnaise increased from $2.75 to $3.25, the consumer increased consumption from 36 six-packs to 44. What is the cross-price elasticity of demand for these two products? What does the calculated elasticity imply about the relationship between peanut butter and mayonnaise for this consumer?


1
Expert's answer
2021-03-23T11:50:09-0400

Solution:

Cross elasticity is used to determine the relationship between two products. It measures the sensitivity of quantity demand change of product X to a change in the price of product Y.

The cross-price elasticity of demand = "\frac{Percentage\; change\; in \;quantity\; demanded}{Percentage \;change\; in \;price}"


Using the point elasticity formula:

Percentage change in quantity demanded = "\frac{(44 - 36) }{36} \times 100 = 22.22\%"


Percentage change in price of mayonnaise = "\frac{(3.25 - 2.75) }{2.75} \times 100 = 18.18\%"


Cross-price elasticity of demand ="\frac{22.22\%}{18.18\%} = 1.22"


Cross-price elasticity of demand = 1.22 > 0


This means that mayonnaise and peanut butter are directly related and are substitutes.